### v1

# From Reliability to Security: How RIS-Assisted Adaptive SM and SSK Enhances Wireless Systems

Abstract
This paper proposes two novel wireless transmission schemes, namely reconfigurable intelligent surface (RIS)-assisted received adaptive spatial modulation (RASM) scheme and RIS-assisted received adaptive space shift keying (RASSK) scheme, designed to enhance spectral efficiency (SE) and physical layer security (PLS). In both proposed schemes, transmitting bits are dynamically mapped at receive antennas by leveraging the characteristics of the RIS in each time slot, which enables the enhancement of signal-to-noise ratio (SNR) at specific selected antennas with near few power, thus leading a reliable and green wireless communication. This adaptive approach facilitates the conveyance of extra bits to the receiver, which means it needs less cost of radio-frequency chains at transmitter while improving SE. Besides, the proposed schemes offer an inherent PLS security advantage, as the eavesdropper is unable to completely detect signals reflected from the RIS. To comprehensively evaluate the performance of the proposed RASM and RASSK schemes, this paper presents a detailed analytical performance of their spectral efficiency, detection complexity, bit error rate, and secrecy rate, which are accompanied by insightful findings and conclusions. Simulation and analytical results demonstrate the superiority of the proposed schemes, showcasing their improved error performance and robustness against wiretapping, while also highlighting the potential of the RASM and RASSK schemes for future wireless applications.
Index Terms:
RIS, IM, spectral efficiency, PLS
I Introduction
I-A Index Modulation
With the exponential growth of devices accessing wireless networks, communication throughput and data rates have surged recently. In response, novel techniques have emerged, with index modulation (IM) standing out. Numerous variants exist in the IM area. Prominent multi-antenna techniques include spatial modulation (SM), space shift keying (SSK), generalized space shift keying (GSSK), and others. These traditional schemes map antenna indices to convey information: single-antenna activation in SM/SSK and multi-antenna combinations in generalized schemes, enabling extra bits and enhancing spectral efficiency (SE) [1]. In IM schemes, constellation symbols modulated by $M$-ary phase shift keying (PSK) or quadrature amplitude modulation (QAM) are transmitted, while additional bits are encoded via antenna indices, each representing distinct transmitted bits. Unlike SM, SSK omits $M$-ary symbols, reducing digital modulation complexity at transmitter and receiver. As noted in [2], SSK exhibits slightly better bit-error-rate (BER) performance than SM with perfect channel state information (CSI) at the receiver. However, conventional SSK [2] demands numerous antennas for higher SE, raising costs and degrading BER. GSSK, introduced in [3], alleviates antenna costs by activating a fixed number, but yields inferior BER compared to SM and SSK. Thus, another groundbreaking IM approach is called addaptive spatial modulation (ASM), which provides flexibility through variable activated transmit antennas forming diverse antenna combinations (ACs), achieving superior SE over traditional SM, SSK, etc. [4]. By minimizing transmit antennas or ACs, ASM’s BER outperforms other IM schemes at equivalent SE. ASM incorporates constellation symbols, increasing complexity relative to adaptive space shift keying (ASSK), which excludes them. To lower the processing complexity at transmitter, this paper also explores ASSK, based on SSK principles where antennas emit pure power, simplifying modulation. As a trade-off, ASM and ASSK enhance SE beyond traditional IM schemes but introduce challenges, such as elevated radio frequency (RF) chain costs and physical layer security (PLS) concerns. To address these, a novel technique is applied to ASM and ASSK for system improvement.
I-B Reconfigurable Intelligent Surface
As a promising technique garnering considerable attention in recent years, reconfigurable intelligent surfaces (RIS) holds great potential for application in future beyond 5th generation (B5G) and 6th generation (6G) wireless communication systems. The RIS comprises a chip, referred to as the intelligent unit, which is connected to the transmitter. This intelligent unit receives information about phase adjustments and subsequently controls the angles of the reflection elements (REs) accordingly [5]. The REs, as fabricated from the special materials that can reflect electromagnetic waves [6], are arranged in a two-dimensional (2D) planar array [7] and can be individually controlled by the intelligent unit. According to it, a deployment feature in the RIS is provided, that is, an extension in coverage of wireless communication through altering the communication delivery into the desired directions [8]. Unlike the traditional relays, the RF chains and amplifiers are not required in the RIS, which can better improve the power utilization in communication. Therefore, through the RIS, the phase shift of transmitted signals can be adjusted with near few powers and the signal-to-noise ratio (SNR) can be further enlarged at receiver, which is named the passive beamforming [9]. By possessing these advantages, the RIS can achieve the IM scheme at receive antennas by adjusting the phase to achieve the passive beamforming at one or more specific antennas, in which it can more effectively address the aforementioned issues of the ASM and ASSK schemes mentioned above and provides a viable consideration of system design for the proposed scheme, as illustrated as follows.
I-C RIS-assisted IM schemes
Given extensive research on reconfigurable intelligent surfaces (RIS), IM is increasingly integrated into its applications. First, RIS acts as a passive relay to enhance IM communication quality. In [10]-[11], an RIS-assisted IM scheme is proposed where SSK operates at transmit antennas, improving BER performance via RIS assistance, as termed TSSK herein for convenience. RIS-assisted SM schemes are explored in various state-of-the-art works [12]-[16], with SM at transmit antennas and RIS boosting receiver SNR, sa called the TSM in this paper. To reduce antennas, RIS-assisted GSSK is introduced in [12], adhering to GSSK activation principles at transmit antennas, though with inferior BER compared to TSSK and TSM. Recent advancements include diverse RIS-assisted IM variants operable not only at transmitters but also at receivers via antenna selection, referred to as received IM herein. As a trade-off in RIS-assisted systems, received IM offers advantages like superior BER, reduced transmitter complexity, and potential gains in SE over related schemes [17]-[21]. In [17], RIS-assisted RSM and RSSK are first proposed with antenna selection at receive antennas, yielding good BER but omitting transmitter-RIS channel due to AP design, with room for SE improvement. Subsequently, RIS-assisted RGSM and RGSSK schemes emerge in [18]-[19], enhancing SE via novel low-complexity detectors, albeit with poorer BER than RSM and RSSK. All these schemes employ fixed selected or activated antennas for information mapping, limiting flexible SE enhancement or dynamic antenna adjustment. To overcome this, we propose two novel schemes with a flexible adaptive antenna selection algorithm, enabling dynamic resource configuration for improved SE adaptability and surpassing fixed-antenna constraints. The RIS-assisted transmitted ASM scheme is termed TASM, with ASM at the transmitter.
I-D Contributions
The existing RIS-assisted IM schemes still face several critical challenges, e.g., contradicting with the low-cost nature of wireless communications to multiple RF chains, limit SE adaptability in dynamic channels by fixing antenna activation patterns, low PLS in wireless applications, etc. To address these issues, we propose two novel RIS-assisted received IM schemes, termed the the RIS-assisted received ASM (RASM) and RIS-assisted RASSK schemes, which dynamically map information bits to receive ACs instead of transmit antennas. Note that prior works on RIS-assisted IM schemes primarily focus on BER or data rate, overlooking the potential of RIS to enhance PLS in this domain. By leveraging RIS’s passive beamforming to enhance SNR at selected receive ACs, our schemes achieve higher SE, while the unique design in reflections inherently enhances the physical layer security against eavesdropping. Overall, the main contributions of this work can be briefly summarized as follows:
  • This paper proposes two novel RIS-assisted IM schemes, denoted as RASM and RASSK, which enhance performance metrics such as BER, spectral efficiency, secrecy rate (SR), and robustness by first employing adaptive antenna selection at the receive antennas and leveraging the RIS to boost the SNR at the receiver.
  • In our schemes, eavesdroppers face challenges in reconstructing transmitted signal due to the inherent security mechanism embedded in the dynamic receive-side mapping. This paper pioneers the theoretical analysis of the SR for these schemes under a passive eavesdropper, bridging the gap in PLS-based studies for RIS-assisted IM schemes.
  • Accounting for side lobes and inevitable interference, the REs are partitioned into constructive and non-constructive components in our schemes with detailed analytical performance. Also, to further reduce the complexity of the proposed schemes, a generalized pre-defined ACs selection at the receiver is designed.
  • This work presents Monte Carlo simulations of the proposed schemes, evaluating BER, SE, and secrecy rate under perfect and imperfect CSI scenarios with artificial noise (AN), ML and maximal ratio combining (MRC) detectors, yielding valuable insights.
I-E Notation
Notation: Boldface letter is used to denote the matrices. $\left(\cdot\right)^{\mathsf{T}}$ and $\left(\cdot\right)^{\mathsf{H}}$ denote the transpose and complex conjugate transpose operations. $\mathsf{diag}(\cdot)$ represents diagonal matrix operation. $\mathbb{C}^{n\times m}$ is the space of $n\times m$ complex-valued matrices. The real part of a complex variable can be denoted as $\mathfrak{R}\{\cdot\}$. $\mathrm{P_{r}}\left(\cdot\right)$ represents the probability of an event. The exponential function is denoted by both $\exp\!\>(\cdot)$ and $e^{\left(\cdot\right)}$, which, while distinct in notation, convey the identical mathematical concept. $\mathrm{Q(}\cdot)$ means the Q-function. $\left\|\cdot\right\|_{2}$ is the second-norm operation, while $\mathbb{E}(\cdot)$ stands for statistical expectation operation. $\jmath\triangleq\sqrt{-1}$ is the imaginary number. $f\left(\cdot\right)$ and $f\left(\cdot\middle|\cdot\right)$ denote the probability density function (PDF) the conditional PDF. $\mathbb{E}\left(\cdot\right)$ represents the operation of expectation. $\left(\begin{array}[]{c}n\\ k\\ \end{array}\right)$ stands for the binomial coefficient from $n$ choose $k$. $\mathcal{C}\mathcal{N}(\mu,\sigma^{2})$ represents the complex Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. $\mathbb{D}(\cdot)$ stand for variance operations.
Figure 1: System models for the RASM and RASSK schemes.
II System Model
II-A Configuration and Wireless Channel
II-A1 Applied Scenarios
This section illustrates the system models for the proposed RIS-assisted received IM schemes, RASM and RASSK, to enhance wireless networks. As depicted in Fig. 1, we consider a SIMO uplink in an IoT scenario where fixed sensors collect environmental data, e.g., temperature, and transmit it to a base station (BS). The sensor employs a single transmit antenna ($N_{t}=1$), while the BS has $N_{r}$ receive antennas. Due to obstructions, the direct link is blocked, relying solely on the RIS path. The RIS comprises $N$ REs and an intelligent unit linked to the transmitter. For instance, in the RASM scheme, bit sequence 010 maps to an AC selecting the first and fourth receive antennas; the intelligent unit directs the first RE subset to amplify the signal toward the first antenna and the second subset toward the fourth. Note that RASM transmits $M$-ary constellation symbols, whereas RASSK conveys only pure power, necessitating more receive antennas in RASSK. We denote $N_{r}^{1}$ for RASM and $N_{r}^{2}$ for RASSK, with $N_{r}^{1}\ll N_{r}^{2}$.
II-A2 Rayleigh Channel Fading
In both system models, the channel matrix of the channel from the transmitter to the receiver is defined as $\mathbf{H}_{1}\in\mathbb{C}^{N\times 1}$, the other from the RIS to the receiver is represented as $\mathbf{H}_{2}\in\mathbb{C}^{N_{r}\times N}$, which both follow the Rayleigh flat fading with zero mean value and unit variance. The diagonal matrix of adjusted phase in the RIS can be given as $\mathbf{\Phi}=\mathsf{diag}(\exp\!\>\left(\jmath\varphi_{1,m}\right),\exp\!\>\left(\jmath\varphi_{2,m}\right),\cdots,\exp\!\>\left(\jmath\varphi_{N,m}\right))$, with $\varphi_{i,m}$ representing the adjusted phase shift of $i$-th reflecting element and for $m$-th receive antenna. Furthermore, we define the wireless channel from the transmit antenna to the $i$-th reflecting element and the one from the $i$-th reflected to the $m$-th receive antenna as $h_{1}^{i}$ and $h_{2}^{i,m}$ with $i=1,2,\cdots,\mathrm{}N$ and $m=1,2,\cdots,\mathrm{}N_{r}$, respectively. $h_{1}^{i}$ and $h_{2}^{i,m}$ are assumed to be independent and identically distributed complex Gaussian random variables with $\mathcal{C}\mathcal{N}(0,1)$, which can be extended to $h_{1}^{i}=\alpha_{i}e^{-\jmath\theta_{i}}$ and $h_{2}^{i,m}=\beta_{i,m}e^{-\jmath\omega_{i,m}}$, with $\alpha$ and $\beta$ representing the Rayleigh factors in various channels, as well $\theta$ and $\omega$ representing the phase shift of these three channels.
II-A3 Specific Antenna Settings
For convenience, we define the vector of receive antenna to represent the various selected antennas in the $r$-th vector of selected AC as $\boldsymbol{v}_{r}=\left[0,1,\cdots,1\right]^{\mathsf{T}}$, where $\boldsymbol{v}_{r}\in\mathbb{C}^{1\times N_{r}}$, 1 represents the selected antennas, and 0 represents the non-selected ones. For example, based on mapping table in Fig. 1, if the transmitted bits sequence $r_{1}=001$ bits are transmitted, the selected antennas in the second AC are second third ones, of which the vector of receive antenna can be denoted as $\boldsymbol{v}_{2}=\left[0,1,1,0\right]^{\mathrm{T}}$. We can further define $v_{i,n}$ stands for the value of $i$-th element in $\boldsymbol{v}_{r}$ as well the $n$-th receive antenna at receiver. The total number of antennas in each AC can be represented as $N_{a}$, with $1\leq N_{a}\leq N_{r}$, which is not a fixed number in each time slot. For example, $N_{a}=1$ to represent bits sequence ’000’ in time slot 1, and $N_{a}=2$ to represent bits sequence ’001’ or ’010’ in time slot 2, as referring in mapping table in Fig. 1. This further enhances the flexiblity in improving SE and antenna configurations. The full set of ACs is given as $\varLambda(r)$, where $r=1,\mathrm{}2,\mathrm{}\cdots,J$ represents $r$-th selected AC and $J$ represents the total number of ACs. Thus, $J$ can be expressed as $J=\sum_{N_{a}}^{N_{r}}{\left(\begin{array}[]{c}N_{r}\\ N_{a}\\ \end{array}\right)}=2^{N_{r}}-1$. Based on the principle of ASM scheme, although $N_{r}$ receive antennas are equipped at receiver, only $N_{a}$ antennas can be selected to convey transmitted signals.
II-A4 Working Principle
The REs are evenly divided into $N_{a}$ parts, which has $N_{E}=\lfloor\frac{N}{N_{a}}\rfloor$ number of REs, and each part is corresponding for one selected receive antenna to achieve the passive beamforming.  [^1] This even division strategy for RE allocation ensures fair resource distribution without real-time optimization. It enables straightforward passive beamforming by assigning equal RE subsets to boost SNR at each selected antenna. The method reduces computational overhead at the intelligent unit, suiting dynamic IoT scenarios with limited processing power. Besides, each antenna not only can receive the signals reflected from the REs that are corresponding for itself but also from the non-corresponding ones. Specifically, as shown in Fig.1, the first receive antenna can receive the signals reflected from the REs corresponding for itself (as shown in the red dashed line) and can be affected by those for the fourth receive antenna (as shown in blue dashed line) in the meanwhile, which leads to the constructive part and non-constructive part in the signals for $m$-th receive antenna, respectively. Notably, all $N_{r}$ receive antennas remain active during transmission, unlike traditional activation-based IM where unselected antennas are deactivated. Instead, the RIS employs passive beamforming to focus gain through constructive reflections on selected antennas while imposing non-constructive interference on others. The RE allocation strategy involves sequential assignment based on the sorted order of selected receive antenna indices in the AC, pre-agreed upon by transmitter and receiver. As shown in the mapping table in Fig. 1, for bit sequence 010 (mapping to antennas 1 and 4), the first RE part amplifies the signal toward antenna 1, and the second toward antenna 4. Similarly, for 100 (mapping to antennas 1, 2, and 3), the first part targets antenna 1, the second antenna 2, and the third antenna 3. This sequential approach enables straightforward implementation at the intelligent unit, which parses the bit sequence to identify the AC and sorts indices before phase adjustment. It ensures fairness by allocating equal RE resources to each selected antenna, avoiding complex optimization and maintaining low overhead in dynamic environments.
II-A5 Rician Channel Fading
To better model a practical wireless environment for the proposed schemes, the Rician channel fading can also be taken into account. This type of fading is appropriate when there is a strong line-of-sight (LoS) path between the RIS and the receiver. The Rician distribution of $h_{2}^{i,n}$ can be given as
$h_{2}^{i,n}=\sqrt{\frac{K}{K+1}}h_{2,Los}^{i,n}+\sqrt{\frac{1}{K+1}}h_{2,NLos}^{i,n},$ (1)
where $h_{2,Los}^{i,n}$ represents the deterministic LoS component and $h_{2,NLos}^{i,n}$ is a zero-mean complex Gaussian random variable accounting for the non-LoS multi-path components. The Rician factor $K\geqslant 0$ denotes the ratio of the power in the LoS component to that in the scattered components. When $K=0$, the model reduces to Rayleigh fading, while higher $K$ values indicate a stronger LoS presence.
II-B RASM Transmission
Assuming total $b_{\mathrm{RASM}}$ bits are required to transmit, $b_{\mathrm{RASM}}$ bits can be divided into $b_{1}$ and $b_{2}$ bits in the RASM scheme, where $b_{1}$ bits are mapped into the indices of selected antenna combinations at receiver and $b_{2}$ bits are mapped into the ones mapping to $M$-ary PSK/QAM symbol transmitted from the transmitter, as shown in Fig. 1. For the selected AC, each one represents different bit sequence, e.g., the third selected AC indicates the bits 010. Moreover, the last bit in sequence is 1, which is represented by $k$-th constellation symbol $x_{k}$ in $M$-ary PSK/QAM constellation symbol vector as $\boldsymbol{x}_{k}=\left[x_{1},x_{2},\cdots,x_{M}\right]^{\mathsf{T}}$ as shown in Fig. 1. The expression for the received signal in the RASM scheme can be written as
$\boldsymbol{Y}=\mathbf{H}_{2}\mathbf{\Phi}_{r}\mathbf{H}_{1}x_{k}+\mathbf{N},$ (2)
where $\mathbf{\Phi}_{r}\in\mathbb{C}^{N\times N}$ represents the diagonal matrix of adjusted phase for $r$-th AC is the vector of Gaussian white noise. However, to better show the differences between various indices of ACs, we give the received signal expression at $l$-th receive antenna in $r$-th selected AC without Gaussian white noise $\tau_{l}$, which is expressed as
$y^{\prime}_{l}=\underset{\mathrm{constructive}\,\,part}{\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{N_{E}}{h_{2}^{i,l}\Phi_{i,l}h_{1}^{i}}x_{k}}}+\underset{\mathrm{non}-\mathrm{constructive}\,\,part}{\underbrace{\sum_{q=1,\mathrm{ }q\neq l}^{N_{a}}{\sum_{i=i^{\prime}}^{qN_{E}}{h_{2}^{i,l}\Phi_{i,l}h_{1}^{i}}x_{k}}}},$ (3)
with $i^{\prime}=\left(q-1\right)N_{E}+1$, as extended as  [^2] We assume that each part of REs can perfectly achieve the passive beamforming for specific $l$-th selected antenna as $\varphi_{i}=\omega_{i,l}+\theta_{i}$ [9]-[12].
$y^{\prime}_{l}=\underset{\mathrm{non-constructive}\,\,part}{\underbrace{\sum_{q=1,\mathrm{ }q\neq z}^{N_{a}}{\sum_{i=i^{\prime}}^{qN_{E}}{\beta_{i,l}\alpha_{i}e^{\jmath\varPsi_{i,l}}x_{k}}}}}+\underset{\mathrm{constructive}\,\,part}{\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{lN_{E}}{\beta_{i,l}\alpha_{i}x_{k}}}},$ (4)
where $\Phi_{i,l}$ can be simplified as $\exp\!\>\left(\jmath\varPsi_{i,l}\right)$ representing the adjust phase of the $i$-th RE and for the $l$-th selected receive antenna in selected AC, $\varPsi_{i,l}=\varphi_{i}-\omega_{i,l}-\theta_{i}$, $\mathrm{}l\in\left\{1,\mathrm{}\cdots,\mathrm{}N_{a}\right\}$. In the meanwhile, the complete received signal at $l$-th selected antenna is given as $y_{l}=y^{\prime}_{l}+\tau_{l}$ with $\tau_{l}$ representing the Gaussian white noise at $l$-th selected receive antenna in selected AC with $\mathcal{C}\mathcal{N}\left(0,N_{0}\right)$ distribution. In Eq. (3) and (4), note that regarding the beam from non-constructive parts and beam narrowness in dense arrays, e.g., $\frac{\lambda}{2}$ spacing, the model incorporates non-constructive reflections, i.e., side lobes and interference, avoiding the assumption of zero interference. And the received signal of $u$-th unselected received antenna can be given as
$y_{u}=\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,u}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}x_{k}}}+\tau_{u},$ (5)
with which the SNR can be obtained as
$\gamma_{u}=\frac{\left\|\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,u}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}x_{k}}}\right\|^{2}}{N_{0}}.$ (6)
According to the characteristic of the RIS, when $\varphi_{i}=\omega_{i,l}+\theta_{i}$, the SNR at $l$-th selected antenna can be maximized, which is given as
$\left\{\gamma_{l}\right\}_{max}=\frac{\left|\sum_{i=\left(l-1\right)N_{E}+1}^{lN_{E}}{\beta_{i,l}\alpha_{i}x_{k}}\right|^{2}}{N_{0}}$
$+\frac{\left|\sum_{q=1,\mathrm{ }q\neq l}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,l}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}x_{k}}}\right|^{2}}{N_{0}}.$ (7)
Applying the ML detector, the expression of ML detection in the RASM scheme can be given as follows:
$\left\{\hat{r},\hat{k}\right\}=\mathrm{arg}\min_{r,k}\!\>\left\|\boldsymbol{Y}-\mathbf{G}_{r,k}\right\|_{2}^{2},$ (8)
where $\hat{r}$ and $\hat{k}$ represent the estimated $r$-th AC and $k$-th $M$-ary PSK/QAM symbol, and $\mathbf{G}_{r,k}=\mathbf{H}_{2}\mathbf{\Phi}_{r}\mathbf{H}_{1}x_{k}$. Here, $\mathbf{G}_{r,k}$ can be rewritten as $\mathbf{G}_{r,k}=\left\{G_{n}\right\}_{n=1}^{N_{r}}$, where $G_{n}$ is
$G_{n}=\begin{cases}\sum_{i=\left(l-1\right)N_{E}+1}^{lN_{E}}{h_{2}^{i,n}h_{1}^{i}}x_{k}+G_{n}^{\prime},\,v_{r,n}=1\\ \\ \sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,u}\alpha_{i}e^{\jmath\varPsi_{i,l}}x_{k}}},\,v_{r,n}=0\end{cases}$ (9)
with $G_{n}^{\prime}=\sum_{q=1,q\neq l}^{N_{a}}{\sum_{i=\varsigma}^{qN_{E}}{h_{2}^{i,n}\Phi_{i,l}h_{1}^{i}}x_{k}}$.
II-C RASSK Transmission
In the RASSK scheme, $b_{\mathrm{RASSK}}$ bits are transmitted to the receiver, which are all mapped into the indices of ACs by the receiver. Similar to the above steps, the received signal expression of the RASSK scheme without the constellation symbols can be given as
$\boldsymbol{Y}=E_{s}\mathbf{H}_{2}\mathbf{\Phi}_{r}\mathbf{H}_{1}+\mathbf{N},$ (10)
where $E_{s}$ represents the transmitted energy. The received signal expression without Gaussian white noise at $l$-th receive antenna for the selected AC in the RASSK scheme is
$y^{\prime}_{l}=$ $\underset{\mathrm{constructive}\,\,part}{\underbrace{E_{s}\sum_{i=\left(l-1\right)N_{E}+1}^{N_{E}}{\beta_{i,l}\alpha_{i}}}}+$
$\underset{\mathrm{non}-\mathrm{constructive}\,\,part}{\underbrace{E_{s}\sum_{q=1,\mathrm{ }q\neq l}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,l}\alpha_{i}e^{\jmath\varPsi_{i,l}}}}}},$ (11)
which, including Gaussian white noise, can be expressed as $\tau_{l}$ as $y_{l}=y^{\prime}_{l}+\tau_{l}$. In the meanwhile, the received signal of $u$-th unselected received antenna in the RASSK scheme is
$y_{u}$ $=E_{s}\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{h_{2}^{i,u}\Phi_{i,l}h_{1}^{i}}}+\tau_{u}$
$=E_{s}\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,u}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}}}+\tau_{u}.$ (12)
Thus, the received SNR at $u$-th unselected antenna is:
$\gamma_{u}=\frac{E_{s}\left\|\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,u}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}}}\right\|^{2}}{N_{0}},$ (13)
and then it can be maximized through $\varphi_{i}=\omega_{i,u}+\theta_{i}$ at $l$-th selected antenna can be given as
$\left\{\gamma_{l}\right\}_{\text{max}}$ $=E_{s}\frac{\left|\sum_{i=(l-1)N_{E}+1}^{lN_{E}}\beta_{i,l}\alpha_{i}\right|^{2}}{N_{0}}$
$\quad+E_{s}\frac{\left|\sum_{q=1,q\neq l}^{N_{a}}\sum_{i=(q-1)N_{E}+1}^{qN_{E}}\beta_{i,l}\alpha_{i}e^{\jmath\varPsi_{i,l}}\right|^{2}}{N_{0}}.$ (14)
The ML detector of the RASSK scheme can be given a
$\left\{\hat{r}\right\}=\mathrm{arg}\min_{r}\!\>\left\|\boldsymbol{Y}-\mathbf{G}_{r}\right\|_{2}^{2}$ (15)
with $\mathbf{G}_{r}=E_{s}\mathbf{H}_{2}\mathbf{\Phi}_{r}\mathbf{H}_{1}$. Meanwhile, $\mathbf{G}_{r}$ can be rewritten as $\mathbf{G}_{r}=\left\{G_{n}\right\}_{n=1}^{N_{r}}$, where $G_{n}$ can be expressed as
[…]
with $G_{n}^{\prime}=E_{s}\sum_{q=1,q\neq l}^{N_{a}}{\sum_{i=\varsigma}^{qN_{E}}{h_{2}^{i,n}\Phi_{i,l}h_{1}^{i}}}$. Similar to example of RASM in Fig. 1, transmitted bit sequence 010 is mapped into 1st and 4th receive antennas, then successfully being detected at receiver.
III Generalized Pre-Defined ACs Selection
Due to the obstruction of the direct link from the RIS to the transmitter and phase adjustment is only optimized for the receiver, obtaining perfect instantaneous CSI from the receiver to the RIS and from the RIS to the transmitter is extremely challenging at the transmitter. Meanwhile, estimating instantaneous CSI for each time slot is not high-efficient due to its high computational complexity and additional spectrum resource requirements, which are especially not practical in future wireless networks.
Since the number of available ACs must be a power of 2 and is more than what is required, e.g., 3 receive antennas lead to 7 ACs but 4 are available to represent the transmitted bit sequences, we need the generalized pre-defined ACs selection to eliminate the unnecessary ACs. Before the transmission, the transmitter and receiver have an agreement that only $D$ ACs can be the indices to convey the bits sequence, where $D=\lfloor\log_{2}\left(N_{r}-1\right)\rfloor$ and $D$ ACs are selected by the designed generalized pre-defined ACs selection. These $D$ ACs are selected through the specific order. [^3] Our aim in this paper is primarily establishing and showing a fundamental system designs and comprehensive performance analysis. The discussion for this extension are out of scope and space limitations, as the potential investigation in future works. In our proposed scheme, we eliminate the AC with the maximum number of antennas. For instance, when $N_{t}=4$, we will remove the ACs with $N_{a}=4$. Then, we sort the remaining ACs in ascending order based on $N_{a}$. The first $D$ ACs are selected for bit mapping. By doing so, we can not only remove the ACs with the maximum number of $N_{a}$ but also minimize the involvement of ACs with higher $N_{a}$ in subsequent signal processing stages, which can further reduce the effects from the non-constructive part. For example, we have $N_{t}=4$ and select 8 ACs based on this order from the 15 available ACs through the aforementioned steps. In this way, the AC with $N_{a}=4$ and most part of ACs with $N_{a}=3$ are eliminated. After this, the AC set can be given as $\varLambda(r)$ with $r=1,2,\cdots,D$. The designed selection can significantly reduce the detection complexity, and a detailed analysis of its detection complexity is provided in sub-section B of Sec. IV. The generalized pre-defined ACs selection works at the intelligent unit of the RIS after the information of selected ACs inputting. Then, the intelligent unit controls various parts of REs to amplify the transmitted signal for the specific receive antenna/antennas.
IV Performance Analysis
In this section, we discuss the performance analysis of the RASM and RASSK schemes in terms of SE, BER, and detection complexity. Firstly, the SE performance of the proposed schemes is presented. Notably, to better demonstrate and understand the SE performance, we utilize the bit per channel use (bpcu) of various IM schemes to represent and discuss the SE performance. Secondly, we derive and analyze the upper bound of the average BER (ABER) to represent the BER performance of both proposed schemes, assuming Rayleigh channel fading and the use of the ML detector. Furthermore, the complexity of ML detector in the proposed schemes and other related schemes is discussed.
IV-A SE Performance
For each time slot, in the RASM scheme, the SE of $b_{1}$ bits mapped into ACs can be expressed as
$\mathsf{SE_{1}}=\lfloor\log_{2}\!\>J\rfloor=\lfloor\log_{2}\!\>\left(2^{N_{r}}-1\right)\rfloor,$ (17)
and the SE of $b_{2}$ bits mapped into constellation symbols can be given as $\mathsf{SE_{2}}=\log_{2}\!\>M$. Thus, the total SE of the RASM can be obtained by $SE_{1}+SE_{2}$ as
$\mathsf{SE_{RASM}}=\lfloor\log_{2}\!\>\left(2^{N_{r}}-1\right)\rfloor+\log_{2}\!\>M.$ (18)
Meanwhile, the SE of the RASSK scheme can be given as
$\mathsf{SE_{RASSK}}=\lfloor\log_{2}\!\>\left(2^{N_{r}}-1\right)\rfloor.$ (19)
In order to show the superiority in the SE of the RASM and RASSK schemes, we also provide the SE expression of the RSM, RSSK schemes, which are given respectively as follows:
$\mathsf{SE_{RSM}}=\log_{2}\!\>N_{r}+\log_{2}\!\>M,$ (20)
$\mathsf{SE_{RSSK}}=\log_{2}\!\>N_{r}.$ (21)
Furthermore, the RGSM and RGSSK schemes, as the schemes with higher SE than that of the RSM and RSSK schemes, have a fixed number of selected antennas as $N_{s}$ in each time slot to convey signals, of which the SE can be given as
$\mathsf{SE_{RGSM}}=\bigg\lfloor\log_{2}\!\>\left(\begin{array}[]{c}N_{s}\\ N_{r}\\ \end{array}\right)\bigg\rfloor+\log_{2}\!\>M,$ (24)
$\mathsf{SE_{RGSSK}}=\bigg\lfloor\log_{2}\!\>\left(\begin{array}[]{c}N_{s}\\ N_{r}\\ \end{array}\right)\bigg\rfloor.$ (27)
Remark 1
Comparing Eq. (18) and Eq. (19), we can easily find that $\mathsf{SE_{RASM}}>\mathsf{SE_{RASSK}}$ with the same $N_{r}$, and the SE of both schemes increases with the increase of $N_{r}$. In order to conduct more impartial SE comparisons, by comparing Eq. (18), Eq. (20), Eq. (24) and Eq. (19), Eq. (21), Eq. (27) respectively, we can draw an insightful conclusion as $\mathsf{SE_{RASM}}\geq\mathsf{SE_{RGSM}}>\mathsf{SE_{RSM}}$ and $\mathsf{SE_{RASSK}}\geq\mathsf{SE_{RGSSK}}>\mathsf{SE_{RSSK}}$ with same $N_{r}$ and $M$, where the equivalence sign can be applicable when $N_{r}$ is large enough and $N_{s}>2$. However, the BER performance of the RGSM and RGSSK schemes has significant deterioration when $N_{s}>2$, but the RASM and RASSK schemes still exhibit better performance while remaining the same $N_{r}$.
IV-B BER Performance
This subsection provides the analytical performance in ABER of both RASM and RASSK schemes, where some insights are also given with $x_{k}=x_{\hat{k}}=\sqrt{E_{s}}$ and Rician channel fading with $K=0$ is considered.
IV-B1 ABER for the RASM Scheme
The analytical ABER performance of the RASM scheme is given. According to Eq. (8) of ML detector, the instantaneous pairwise error probability (PEP) can be expressed by giving channel $\mathrm{h}_{1}^{i}$ and $\mathrm{h}_{2}^{i,l}$ as
$\mathrm{P_{r}}\left(\left\{r,k\right\}\rightarrow\{\hat{r},\hat{k}\}\mid h_{1}^{i},h_{2}^{i,m}\right)=\mathrm{Pr}\left(\!\>\left\|\boldsymbol{Y}-\mathbf{G}_{r,k}\right\|_{2}^{2}\right)$
$=\mathrm{Pr}\!\>\left(\sum_{n=1}^{N_{r}}{\left\|y_{n}-G_{n}\right\|_{2}^{2}}>\sum_{n=1}^{N_{r}}{\left\|y_{n}-\hat{G}_{n}\right\|_{2}^{2}}\right),$ (28)
where $\hat{G}_{n}$ represents the estimated $G_{n}$ and $y_{n}$ represents the received signal at $n$-th receive antenna. After some algebraic operations, Eq. (28) can be extended and rewritten as
$\mathrm{P_{r}}\left(\left\{r,k\right\}\rightarrow\{\hat{r},\hat{k}\}\mid\mathrm{h}_{1}^{i},\mathrm{h}_{2}^{i,m}\right)=\mathrm{Pr}\!\>\left(\Gamma<0\right),$ (29)
with
$\Gamma=\sum_{n=1}^{N_{r}}{\left\|G_{n}-\hat{G}_{n}\right\|_{2}^{2}}+\sum_{n=1}^{N_{r}}{2\mathfrak{N}\left\{{\tau_{n}}^{*}\left[G_{n}-\hat{G}_{n}\right]\right\}},$ (30)
where $\tau_{l}$$^{*}$ represents the complex conjugation of Gaussian white noise and $\Gamma\sim\mathcal{N}\left(\mu_{\Gamma},\sigma_{\Gamma}^{2}\right)$ with $Z=\sum_{n=1}^{N_{r}}{\left\|G_{n}-\hat{G}_{n}\right\|_{2}^{2}}$ and $\sigma_{\Gamma}^{2}=2N_{0}\sum_{n=1}^{N_{r}}{\left\|G_{n}-\hat{G}_{n}\right\|_{2}^{2}}$. Thus, by applying the Gaussian Q-function as $Q(\frac{-\mu_{\Gamma}}{\sigma_{\Gamma}^{2}})$ the PEP can be further denoted as follows:
$\mathrm{P_{r}}\left(\left\{r,k\right\}\rightarrow\{\hat{r},\hat{k}\}\mid\mathrm{h}^{i},\mathrm{h}_{2}^{i,m}\right)=Q\left(\sqrt{\frac{Z}{2N_{0}}}\right).$ (31)
Considering the alternative form of Q-function, the unconditional PEP, which is averaged over channel coefficients as $\mathrm{P_{r}}\left(\left\{r,k\right\}\rightarrow\{\hat{r},\hat{k}\}\right)=\mathbb{E}_{Z}\left[Q\left(\sqrt{\frac{Z}{2N_{0}}}\right)\right]$, is calculated as
[…]
$=\frac{1}{\pi}\int_{0}^{\frac{\pi}{2}}{M_{Z}\left(\frac{-1}{4\sin^{2}\!\>\tau N_{0}}\right)d\tau},$ (32)
which applies the moment generating function (MGF) of $Z$, given as $M_{Z}\left(t\right)=\int_{0}^{\infty}{e^{tZ}dZ}$, with $t=\frac{-1}{4\sin^{2}\!\>\tau N_{0}}$. Actually, $Z$ is the squared Euclidean distance (SED) between the symbols $G_{r,k}$ and $G_{\hat{r},\hat{k}}$. Here, the MGF of Eq. (32) can be derived by considering the general quadratic form of correlated Gaussian random variables and counts on erroneous or correct detection of the $r$-th and $\hat{r}-$th AC indices. Note that the error in the detection of AC index is able to affect the accuracy of constellation symbol detection. Thus, we separate $Z$ into two categories, which are given as: $Z_{1}=\left\{Z\mid r\neq\hat{r}\right\}$ and $Z_{2}=\left\{Z\mid r=\hat{r}\right\}$.
1) First case $Z_{1}$: $\left\{Z\mid r\neq\hat{r}\right\}$
In this case, the SED can be given as $Z_{1}=Z_{1}^{1}+Z_{1}^{2}+Z_{1}^{3}$, which are given specifically in Eq. (33), Eq. (34), and Eq. (35) with $\varphi_{i}-\omega_{i,l}-\theta_{i}=\varPsi_{i,l}$, as shown at the top of next page.
$Z_{1}^{1}=\sum\nolimits_{l\in V_{r}}{\left|\left(\sum_{i=\left(l-1\right)N_{E}+1}^{lN_{E}}{\beta_{i,l}\alpha_{i}}+\sum_{q=1,\mathrm{ }q\neq l}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,l}\alpha_{i}e^{\jmath\varPsi_{i,l}}}}\right)x_{k}-\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,l}\alpha_{i}e^{\jmath\varPsi_{i,l}}x_{\hat{k}}}}\right|}^{2},$ (33)
$Z_{1}^{2}=\sum\nolimits_{{\hat{l}}\in V_{\hat{r}}}{\left|\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,\hat{l}}\alpha_{i}e^{\jmath\varPsi_{i,\hat{l}}}x_{k}}}-\left(\sum_{i=\left(l-1\right)N_{E}+1}^{lN_{E}}{\beta_{i,\hat{l}}\alpha_{i}}+\sum_{q=1,\mathrm{ }q\neq l}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,\hat{l}}\alpha_{i}e^{\jmath\varPsi_{i,\hat{l}}}}}\right)x_{\hat{k}}\right|}^{2},$ (34)
$Z_{1}^{3}=\sum_{n=1,n\neq l,n\neq\hat{l}}^{N_{r}}{\left|\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\left(\beta_{i,n}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}x_{k}-\beta_{i,n}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}x_{\hat{k}}\right)}}\right|^{2}}.$ (35)
Here, $Z_{1}^{1}$, $Z_{1}^{2}$, and $Z_{1}^{3}$ stand for the $m=r$, $m=\hat{r}$, and $m\neq r,\mathrm{}\hat{r}$ respectively. Then, we can denote $Z_{1}^{1}$ and $Z_{1}^{2}$ as
$Z_{1}^{1}=\left|q_{1}\right|^{2}=\left(q_{1}\right)_{\mathfrak{N}}^{2}+\left(q_{1}\right)_{\mathfrak{T}}^{2},$ (36)
$Z_{1}^{2}=\left|q_{2}\right|^{2}=\left(q_{2}\right)_{\mathfrak{N}}^{2}+\left(q_{2}\right)_{\mathfrak{T}}^{2},$ (37)
where $q_{1}$ and $q_{2}$ follow complex Gaussian distribution by applying central limit theorem (CLT).  [^4] Note that $\alpha$ and $\beta$ are the magnitudes of standard complex Gaussian random variables, they follow the Rayleigh distribution with mean value $\frac{\sqrt{\pi}}{2}$ variance $\frac{4-\pi}{4}$, respectively. Moreover, since $\omega_{i}$, $\varphi_{i}$, and $\theta_{i}$ are the phase of standard complex Gaussian random variable and follow the uniform distribution at the range in $(-\pi,\pi)$. The mean vector $\boldsymbol{\mu}_{1}$ and covariance matrix $\boldsymbol{\sigma}_{1}^{2}$ of $\boldsymbol{z}_{1}=\left[\left(q_{1}\right)_{\mathfrak{R}},\left(q_{1}\right)_{\mathfrak{T}},\left(q_{2}\right)_{\mathfrak{R}},\left(q_{2}\right)_{\mathfrak{T}}\mathrm{}\right]$, which are respectively given as follows:
$\boldsymbol{\mu}_{1}=\frac{N\pi}{4}\left[\left(x_{k}\right)_{\mathfrak{N}},\left(x_{k}\right)_{\mathfrak{T}},-\left(x_{k}\right)_{\mathfrak{N}},-\left(x_{k}\right)_{\mathfrak{T}}\right]^{\mathsf{T}},$ (38)
$\boldsymbol{\sigma}_{1}^{2}=\begin{bmatrix}\vskip 2.15277pt\sigma_{1}^{2}&\sigma_{1,2}^{2}&\sigma_{1,3}^{2}&\sigma_{1,4}^{2}\\ \vskip 2.15277pt\sigma_{1,2}^{2}&\sigma_{2}^{2}&\sigma_{2,3}^{2}&\sigma_{2,4}^{2}\\ \vskip 2.15277pt\sigma_{1,3}^{2}&\sigma_{2,3}^{2}&\sigma_{3}^{2}&\sigma_{3,4}^{2}\\ \vskip 2.15277pt\sigma_{1,4}^{2}&\sigma_{2,4}^{2}&\sigma_{3,4}^{2}&\sigma_{4}^{2}\\ \end{bmatrix}$, (39)
where
$\sigma_{1}^{2}=N_{a}\left[\left(N_{a}-\frac{\pi^{2}}{16}\right)N_{E}\left(x_{k}\right)_{\mathfrak{R}}^{2}+\frac{N\left|x_{\hat{k}}\right|^{2}\mathrm{ }}{2}\right],$
$\sigma_{2}^{2}=N_{a}\left[\left(N_{a}-\frac{\pi^{2}}{16}\right)N_{E}\left(x_{k}\right)_{\mathfrak{T}}^{2}+\frac{N\left|x_{\hat{k}}\right|^{2}\mathrm{ }}{2}\right],$
$\sigma_{3}^{2}=N_{a}\left[\left(N_{a}-\frac{\pi^{2}}{16}\right)N_{E}\left(x_{\hat{k}}\right)_{\mathfrak{R}}^{2}+\frac{N\left|x_{k}\right|^{2}\mathrm{ }}{2}\right],$
$\sigma_{4}^{2}=N_{a}\left[\left(N_{a}-\frac{\pi^{2}}{16}\right)N_{E}\left(x_{\hat{k}}\right)_{\mathfrak{T}}^{2}+\frac{N\left|x_{k}\right|^{2}\mathrm{ }}{2}\right].$ (40)
By applying the property of covariance between two random variables $X$ and $Y$ as
$2\mathrm{Cov}\left(X,Y\right)=\mathbb{D}\left[X+Y\right]-\mathbb{D}\left[X\right]-\mathbb{D}\left[Y\right],$ (41)
we can easily obtain the $\sigma_{1,2}^{2}$, $\sigma_{3,4}^{2}$, $\sigma_{1,4}^{2}$, $\sigma_{2,3}^{2}$, and $\sigma_{2,4}^{2}$. The MGF of the generalized non-central chi-square distribution is given as follows [11]-[12]:
$M_{X}\left(t\mid\boldsymbol{\mu},\boldsymbol{\sigma}^{2}\right)$ $=\left[\det\left(\mathbf{E}-2t\boldsymbol{\sigma}^{2}\right)\right]^{-\frac{1}{2}}\times$
$\exp\left\{-\frac{1}{2}\boldsymbol{\mu}^{\mathsf{T}}\left[\mathbf{E}-\left(\mathbf{E}-2t\boldsymbol{\sigma}^{2}\right)^{-1}\right]{\boldsymbol{\sigma}}^{-1}\boldsymbol{\mu}\right\},$ (42)
where $X=\sum_{f=1}^{f}{X_{f}^{2}}$, is the unit matrix, $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}^{2}$ represent the mean vector and covariance matrix of $\left[X_{1},X_{2},\cdots,X_{f}\right]^{\mathsf{T}}$. Then, substituting Eq. (38) and Eq. (39) into Eq. (42), we can yield the MGF of $Z_{1}^{1}+Z_{1}^{2}$, given as $M_{Z_{1}^{1}+Z_{1}^{2}}\left(t\mid\boldsymbol{\mu}_{1},\boldsymbol{\sigma}_{1}^{2}\right)$. Nevertheless, the MGF of $Z_{1}^{3}$ is still needed to be derived. In Eq. (35), we define that
$\Gamma=$ $\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,n}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}x_{k}}}$
$-\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,n}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}x_{\hat{k}}}}$, (43)
which has variance $\sigma_{\Gamma}^{2}=\frac{N\left(\left|x_{k}\right|^{2}+\left|x_{\hat{k}}\right|^{2}\right)}{2}$ and zero mean value. For Eq. (35), $Z_{1}^{3}$ follows the generalized central chi-square distribution with $2\left(D-2\right)N_{a}$ degree of freedom and has MGF:
$M_{Z_{1}^{3}}\left(t\right)=\left[1-tN\left(\left|x_{k}\right|^{2}+\left|x_{\hat{k}}\right|^{2}\right)\right]^{\frac{-N_{l}}{2}},$ (44)
where $N_{l}$ represents the number of same antennas in the selected AC and estimated AC. Finally, the MGF of $Z_{1}$ can be obtained as
$M_{Z_{1}}\left(t\right)=M_{Z_{1}^{1}+Z_{1}^{2}}\left(t\mid\boldsymbol{\mu}_{1},\boldsymbol{\sigma}_{1}^{2}\right)M_{Z_{1}^{3}}\left(t\right).$ (45)
And the unconditional PEP of case 1 can be given by substituting Eq. (45) into Eq. (32) as
[…]
$=\frac{1}{\pi}\int_{0}^{\frac{\pi}{2}}{M_{Z_{1}^{1}+Z_{1}^{2}}\left(t\mid\boldsymbol{\mu}_{1},\boldsymbol{\sigma}_{1}^{2}\right)M_{Z_{1}^{3}}\left(t\right)d\tau}.$ (46)
2) Second case $Z_{2}$: $\left\{Z\mid r=\hat{r}\right\}$:
In this case, we also can divide $Z_{2}$ into $Z_{2}^{1}+Z_{2}^{2}$ with
$Z_{2}^{1}=$ $\left|x_{k}-x_{\hat{k}}\right|^{2}\sum_{l\in V}{\left|\sum_{q=1,q\neq l}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,l}}}\alpha_{i}e^{\jmath\Psi_{i,l}}\right|^{2}}$
$+\left|x_{k}-x_{\hat{k}}\right|^{2}\sum_{l\in V_{r}}{\left|\sum_{i=\left(l-1\right)N_{E}+1}^{N_{E}}{\beta_{i,l}}\alpha_{i}\right|^{2}},$ (47)
$Z_{2}^{2}=\left|x_{k}-x_{\hat{k}}\right|^{2}\sum_{n=1,n\neq l}^{N_{r}}{\left|\sum_{q=1}^{N_{a}}{\sum_{i=i^{\prime}}^{qN_{E}}{\beta_{i,n}\alpha_{i}e^{\jmath\left(\varphi_{i}-\omega_{i,l}-\theta_{i}\right)}}}\right|^{2}},$ (48)
where $Z_{2}^{1}$ and $Z_{2}^{2}$ stand for the situations that $m=r$ and $m\neq r$, respectively. Firstly, by approximating $\beta_{i,l}\alpha_{i}$ as a Gaussian random variable with CLT, the mean value and variance of $Z_{2}^{1}$ can be obtained as $\mu_{Z_{2}^{1}}=\frac{N_{a}N_{E}\pi\left|x_{k}-x_{\hat{k}}\right|}{4}$ and $\sigma_{Z_{2}^{1}}^{2}=\frac{\left|x_{k}-x_{\hat{k}}\right|^{2}N_{a}N_{E}\left(32-\pi^{2}\right)}{16}$. Then, by substituting $\mu_{Z_{2}^{1}}$ and $\sigma_{Z_{2}^{1}}^{2}$ into Eq. (42), the MGF of $Z_{2}^{1}$ can be obtained as
$M_{Z_{2}^{1}}\left(t\mid\mu_{Z_{2}^{1}},\sigma_{Z_{2}^{1}}^{2}\right)=\left(1-2\sigma_{Z_{2}^{1}}^{2}t\right)^{-\frac{1}{2}}e^{\frac{{t\mu_{Z_{2}^{1}}}^{2}}{1-2\sigma_{Z_{2}^{1}}^{2}t}}.$ (49)
Secondly, $\Gamma_{2}$ can be approximately considered as a Gaussian random variable applying CLT:
$\Gamma_{2}=\sum_{q=1}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\beta_{i,n}\alpha_{i}e^{\jmath\varPsi_{i,l}}x_{k}}},$ (50)
with zero mean value and variance $\sigma_{\Gamma_{2}}^{2}=N$. Thus, $Z_{2}^{2}$ follows the generalized non-central chi-square distribution with $(N_{r}-N_{a})$ degree of freedom and has the variance denoted as $\sigma_{Z_{2}^{2}}^{2}=\frac{NN_{a}\left|x_{k}-x_{\hat{k}}\right|^{2}}{2}$. Giving $\sigma_{Z_{2}^{2}}^{2}$ and following Eq. (45), the MGF of $Z_{2}^{2}$ yields as
$M_{Z_{2}^{2}}\left(t\right)=\left(1-tN_{a}N\left|x_{k}-x_{\hat{k}}\right|^{2}\right)^{-\frac{N_{r}-N_{a}}{2}}.$ (51)
Thus, the MGF of $Z_{2}$ can be obtained by multiplying Eq. (49) and Eq. (51), which is specifically given as $M_{Z_{2}}\left(t\right)=M_{Z_{2}^{1}}\left(t\right)M_{Z_{2}^{2}}\left(t\right)$, which can be substituted into Eq. (32) to further obtain the unconditional PEP of case 2  [^5] Following the upper bound of Q-function, the PEP can be expressed as $\mathrm{P_{r}}\left(\left\{r,k\right\}\rightarrow\{\hat{r},\hat{k}\}\right)\leq\frac{1}{6}M_{Z}\left(-\frac{1}{N_{0}}\right)+\frac{1}{12}M_{Z}\left(-\frac{1}{2N_{0}}\right)+\frac{1}{4}M_{Z}\left(-\frac{1}{4N_{0}}\right)$ [22]. The closed-form expression can be obtained in this way, which is not extended in this paper. as
$\mathrm{P_{r}}_{Z_{2}}\left(\left\{r,k\right\}\rightarrow\{\hat{r},\hat{k}\}\right)=\frac{1}{\pi}\int_{0}^{\frac{\pi}{2}}{M_{Z_{2}}\left(t\right)d\tau}.$ (52)
Eventually, the union bound of ABER for the RASM scheme can be derived from the values of unconditional PEP by Eq. (28), which is given as
$P_{b}^{RASM}\leq\frac{\sum_{r}{\sum_{\hat{r}}{\sum_{k}{\sum_{\hat{k}}{\mathrm{P}_{\mathrm{r}}\left(\left\{r,k\right\}\rightarrow\left\{\hat{r},\hat{k}\right\}\right)e^{\prime}}}}}}{MD\log_{2}\!\>\left(MD\right)},$ (53)
[…]
IV-B2 ABER for the RASSK Scheme
The analytical ABER performance of the RASSK scheme with the ML detector is given in this sub-section. In light of the steps in ABER derivations of the RASM scheme with the ML detector, the conditional PEP of it can be given:
$\mathrm{P_{r}}\left(r\rightarrow\hat{r}\middle|\mathrm{h}_{1}^{i},\mathrm{ h}_{2}^{i,m}\right)\overset{\mathrm{def}}{=}Q\left(\sqrt{\frac{B}{2N_{0}}}\right),$ (54)
where $B=\sum_{n=1}^{N_{r}}{\left\|G_{n}-G_{\hat{n}}\right\|}_{2}^{2}$ Then, by averaging over channel coefficients, the average unconditional PEP can be obtained as
$\mathrm{P_{r}}\left(r\rightarrow\hat{r}\right)=\frac{1}{\pi}\int_{0}^{\small{\frac{\pi}{2}}}{M_{B}\left(\frac{-1}{4\sin^{2}\!\>\tau N_{0}}\right)d\tau}.$ (55)
Since the constellation symbols are not transmitted in the RASSK scheme and $x_{k}=x_{\hat{k}}=\sqrt{E_{s}}$., case 1 in the ABER analysis of the RASM scheme with the ML detector is also suitable for the RASSK one where $r\neq\hat{r}$. In other words, the MGF given in Eq. (42) of the RASM scheme in case 1 can be also employed in the RASSK scheme with some modifications in $x_{k}$ and $E_{s}$, i.e., $M_{B}\left(\frac{-1}{4\sin^{2}\!\>\tau N_{0}}\right)\approx M_{Z_{1}}\left(t\right)$. Finally, the upper bound of the RASSK scheme with the ML detector can be obtained similar to
$P_{b}^{RASSK}\leq\frac{\sum_{r}{\sum_{\hat{r}}{\mathrm{P}_{\mathrm{r}}\left(r\rightarrow\hat{r}\right)e\left(r\rightarrow\hat{r}\right)}}}{D\log_{2}\!\>\left(D\right)}.$ (56)
Remark 2
Since the increase of $N$ leads to the increase of $N_{E}$ and probably $N_{a}$, we can find that the results of above MGFs decrease with the arise of $N$, which means that the ABER reduces with the $N$ increase.
IV-C Detection Complexity
In this section, the complexity of the ML detector in the RASM and RASSK schemes are analyzed. We calculate the complexity based on Eq. (8) in the RASM scheme and Eq. (15) in the RASSK scheme, where the complex multiplications (CM) and complex additions (CA) are all required to count [23]-[25]. By doing this, the different steps in Eq. (8) to count the operations of CM and CA in the RASM can be given respectively as follows:
  • $N_{r}(N+1)+N$ operations of CM and $N_{r}(N-1)+N_{r}$ ones of CA are needed in $\boldsymbol{Y}$.
  • Similar to step 1, $\mathbf{G}_{r,k}$ has $N_{r}(N+1)+N$ operations of CM but only $N_{r}(N-1)$ ones of CA.
  • The subtraction operation of $\boldsymbol{Y}-\mathbf{G}_{r,k}$ involves $N_{r}$ ones of CA.
  • In the Frobenius norm, $N_{r}$ CMS and $N_{r}$ CAs are needed for calculate the magnitude for each element, and $N_{r}-1$ CAs are required in summing the magnitudes.
  • In summary, for each set $\left\{r,k\right\}$, total complexity can be given as $2N_{r}(N+2)+N$ of CMs and $2N_{r}(N-1)+4N_{r}-1$ of CAs.
  • Lastly, the total number of all set $\left\{r,k\right\}$ is $D\times M$.
Consequently, the total ML detector’s operations of CM and CA in the RASM can be obtained as
$C_{RASM}=\left[2N_{r}(2N+1+2N_{r})+N-1\right]DM.$ (57)
As following the same steps of the RASM schemes, $\left\|\boldsymbol{Y}-\mathbf{G}_{r}\right\|_{2}^{2}$ term has the operations of CM and CA in the ML detector applied in the RASSK scheme can be derived as $2N_{r}(N+2)+N$ and $2N_{r}(N-1)+4N_{r}-1$, respectively. Differing from the RASM scheme, the total operations of the ML detector in the RASSK scheme can be given as
$C_{RASSK}=\left[2N_{r}(2N+1+2N_{r})+N-1\right]D.$ (58)
Remark 3
In cases where the generalized pre-defined ACs selection is not utilized, $D$ in Eq. (57) and Eq. (58) is replaced into $J$ and $J\gg D$, especially when $N_{r}$ is large. This leads to a substantial increase in detection complexity, thereby highlighting the significant contribution of the designed ACs selection in reducing complexity.
Remark 4
Comparing the Eq. (57) and Eq. (58), we can find the detection complexity of the RASSK schemes is lower than that of the RASM scheme with the same $N_{r}$ and $N$, because the complexity of the detection in constellation symbols is not considered in the RASSK scheme. Also, according to Eq. (57) and Eq. (58), the complexity of both schemes increases with the rise in $N_{r}$ and $N$.
Remark 5
In the RSM and RSSK schemes [16], all receive antennas account for channel fading between the transmitter and the RIS and receive signals reflected from the RIS. The CM and CA operations in these schemes are analogous to those in the RASM and RASSK schemes under the same system model configurations, including $N$ and $N_{r}$, except for SE. Specifically, the CM and CA operations for RASM and RASSK can be expressed similarly using Eq. (57) and Eq. (58), with $D$ replaced by $N_{r}$. This indicates that the complexity of RASM and RASSK approaches that of RSM and RSSK for small $N_{r}$. Nevertheless, the RASM and RASSK schemes achieve higher SE compared to RSM and RSSK with the same $N_{r}$. Thus, as the trad-off options applied future wireless networks, we can find that the proposed schemes can yield acceptable complexity and better SE compared to the RSM and RSSK schemes with small $N_{r}$.
Figure 2: System model of the RASM and RASSK schemes eavesdropper.
V System Model and Secrecy Rate for PLS
V-A System Model with Eavesdropper
Since the adaptive antenna selection works at a specific receiver in both proposed schemes, the eavesdropper cannot completely detect and decode the received signals from the RIS, which further indicates that the proposed schemes offer better secrecy performance in terms of PLS during transmission. A new system model in Fig. 2 includes an additional eavesdropper, referred to as Eve equipped with one antenna,  [^6] We primarily emphasize the superiority of PLS performance in our proposed schemes through comparisons with related benchmarks. Accordingly, Eve is assumed to possess equivalent radio hardware to Bob, e.g., single RF chain, identical noise variance, but lacks Bob’s private seed for RIS reconfiguration. Eve can access statistical channel parameters for transmitter-to-RIS and RIS-to-receiver links, and measure her own instantaneous realizations via public pilots. However, scenarios involving Eve’s physical access to the RIS controller or active jamming, e.g., active eavesdropping or jamming attacks, represent different threat models beyond our scope and are not considered. which aims at eavesdropping the transmitted signals from the RIS to the receiver as Bob. It is worth noting that this system model in Fig. 2 is suitable for both the RASM and RASSK schemes, with the only difference that the constellation symbols are transmitted in the RASM scheme instead of the RASSK one. Besides, Eve is assumed to be passive and placed as close to Bob as possible, which means Eve cannot actively attack or interrupt. Here, we also assume that Eve has perfect knowledge of the CSI from the RIS to the receive antenna and the transmitter to the RIS which is given as $g_{2}\sim\mathcal{C}\mathcal{N}\left(0,1\right)$, following Rayleigh distribution [26]-[28]. Thus, the received signals at Eve can be given as
$y_{e}=\sum_{i=1}^{N}{g_{2}^{i}\Phi_{i}h_{1}^{i}x_{k}}+\tau_{e},$ (59)
where $g_{2}^{i}$ and $h_{1}^{i}$ bit sequence $i$-th RE respectively, $\Phi_{i}$ stands for the adjusted phase from the $i$-th RE, and $n_{e}$ is the white Gaussian noise with variance $N_{0}$. The ML detector at Eve is
$\left\{\hat{k}\right\}=\mathrm{arg}\min_{k}\!\>\left\|y_{e}-\Delta_{e}\right\|_{2}^{2}.$ (60)
where $\Delta_{e}\overset{\mathrm{def}}{=}\sum_{i=1}^{N}{g_{2}^{i}\Phi_{i}h_{1}^{i}x_{k}}$. In this section, we provide a simple analysis in the SR of the RASM and RASSK schemes.
V-B Secrecy Rate Analysis for the RASM Scheme
Referring by [29], the SR of the RASM can be given as
$R_{RASM}=\max\!\>\left\{0,\mathrm{}R_{B}-R_{E}\right\},$ (61)
where $\mathrm{}R_{B}$ and $R_{E}$ represent the data rate of Bob and Eve in bpcu, respectively. Note that each selected AC has the same probability as $\frac{1}{D}$ to be chosen to convey bits sequence in a time slot. Thus, by assuming the data rate for all selected ACs are the same, i.e., $R_{B}^{1}=R_{B}^{2}=\cdots=R_{B}^{D}$ and $R_{E}^{1}=R_{E}^{2}=\cdots=R_{E}^{D}$, we can easily find that $R_{B}=R_{B}^{r}$ and $R_{E}=R_{E}^{r}$, where $R_{B}^{r}$ and $R_{E}^{r}$ represent the Bob’s and Eve’s data rate concerning with $r$-th selected AC. Thus, the data rate of Bob can be further expressed as
$R_{B}=R_{B}^{r}+R_{B}^{k},$ (62)
where $R_{B}^{r}$ and $R_{B}^{k}$ represent the achievable data rate of Bob related to the bits of $r$-th selected AC and $k$-th constellation symbol, respectively. On the contrary, Eve has probability in $\frac{1}{D}$ and $\frac{1}{M}$ to successfully guess and decode the ACs and constellation symbols respectively [29]-[30], so the achievable data rate of it can be given as  [^7] Total number of possible mapping patterns for constellation symbols and ACs that the transmitter can adopt are $M$ and $D$, respectively. The probability that Eve adopts the same mapping pattern as the transmitter is $\frac{1}{M}$ for the constellation symbols and $\frac{1}{D}$ for the ACs. This implies that there is only a $\frac{1}{M}$ chance that Eve can correctly guess the exact mapping pattern used by the transmitter for the constellation symbols, and a $\frac{1}{D}$ chance for the ACs.
$R_{E}=D^{-1}R_{E}^{r}+M^{-1}R_{E}^{k},$ (63)
with $R_{E}^{r}$ and $R_{E}^{k}$ representing the achievable data rate of Bob related to the bits of $r$-th selected AC and $k$-th constellation symbol when Eve successfully decodes both.
According to [29]-[30], the expressions for the achievable data rates of Bob and Eve are given in Eq. (64) and Eq. (65) at the top of the next page, respectively. Detailed proofs can be found in Appendix A.
$R_{B}=\log_{2}\!\>DM-\frac{1}{DM}\sum_{r=1}^{D}{\sum_{k=1}^{M}{\mathbb{E}_{\mathrm{h}_{1}^{i},\mathrm{h}_{2}^{i,n},\tau_{n}}\left\{\log_{2}\!\>\left[1+\sum_{\hat{r}=1,\hat{r}\neq r}^{D}{\sum_{\hat{k}=1,\hat{k}\neq k}^{M}{\exp\left(-\left\|\boldsymbol{Y}-\mathbf{G}_{r,k}\right\|_{2}^{2}+\mathbf{N}\right)}}\right]\right\}}}.$ (64)
$R_{E}=\frac{1}{M}\left\{\log_{2}\!\>M-\frac{1}{M}\sum_{k=1}^{M}{\mathbb{E}_{n_{e},g_{1}^{i},g_{2}^{i}}\left[\log_{2}\!\>\frac{1+\exp\left(-\left\|\tau_{e}\right\|_{2}^{2}\right)}{\sum_{\hat{k}=1}^{M}{\left[1+\exp\left(-\left\|y_{e}-\sum_{i=1}^{N}{g_{2}^{i}\Phi_{i}g_{1}^{i}x_{\hat{k}}}\right\|_{2}^{2}\right)\right]}}\right]}\right\}.$ (65)
V-C Secrecy Rate Analysis for the RASSK Scheme
Similar to Eq. (61), the SR of the RASSK is given as
$R_{RASSK}=\max\!\>\left\{0,\mathrm{}R_{B}-R_{E}\right\},$ (66)
where $R_{B}=R_{B}^{r}$ and $R_{E}=0$ due to the absence of constellation symbols in the RASSK scheme and incapability in detecting the symbols of selected ACs. Furthermore, the mutual information of Bob in the RASSK scheme can be calculated similar to the steps of $\mathfrak{T}_{b}\left(k;\boldsymbol{Y}\middle|\mathrm{h}_{1}^{i},\mathrm{h}_{2}^{i,n}\right)$ with replacing $x_{k}$ into $r$-th AC in $\varLambda(r)$ and some algebraic calculations, which can be obtained as $R_{RASSK}=\max\!\>\left\{0,\mathrm{}R_{B}\right\}$, where $R_{B}=R_{B}^{r}$ and $R_{E}=0$ due to the absence of constellation symbols in the RASSK scheme and incapability in detecting the symbols of selected ACs. Furthermore, the mutual information of Bob in the RASSK scheme can be calculated similar to the steps of $\mathfrak{T}_{b}\left(k;\boldsymbol{Y}\middle|\mathrm{h}_{1}^{i},\mathrm{h}_{2}^{i,n}\right)$ with replacing $x_{k}$ into $r$-th AC in $\varLambda(r)$ and some algebraic calculations. Next, taking the expectation of it over $\mathrm{h}_{1}^{i}$ and $\mathrm{h}_{2}^{i,n}$, $R_{B}$ can be further expressed as
$R_{B}$ $=\log_{2}\!\>D$-
$\frac{1}{D}\sum_{k=1}^{M}{\sum_{\hat{r}=1}^{D}{\mathbb{E}_{\mathrm{h}_{1}^{i},\mathrm{h}_{2}^{i,n},\tau_{n}}\left[\log_{2}\exp\left(\left\|\mathbf{N}\right\|_{2}^{2}-\left\|\boldsymbol{Y}-\mathbf{G}_{r}\right\|\right)\right]}}.$ (67)
Substituting Eq. (67) and $R_{E}=0$ into $R_{RASSK}=\max\!\>\left\{0,\mathrm{}R_{B}\right\}$, we can easily find the final SR through Monte Carlo simulations shown in Sec. VI.
Figure 3: Analytical and simulation BER results of the RASM in (a) and (b), and RASSK schemes in (c).
Figure 4: (a) BER performance and (b) spectral efficiency represented by bpcu of various IM schemes with different $N_{d}$, $N_{r}=16$, $M=2$, and $N=8$.
VI Simulation Results
In this section, the analytical and simulation results are presented. To provide a more comprehensive illustration of the performance in the proposed scheme, a uniform assumption of Rayleigh channel fading is adopted, excepted for Fig. 6, with the transmission antenna $N_{t}=1$ to be considered in all cases of the RASM and RASSK schemes. All of these results are conducted by employing the Monte Carlo simulations, encompassing a substantial number of channel realizations, precisely $3\times 10^{6}$. Similar to [10]-[14], $\bar{\gamma}=\frac{E_{s}}{N_{0}}$ is considered as the average SNR at receiver.
The analytical and simulation results of BER in the RASM scheme are presented in Fig. 3, where the results of the cases with $N_{r}=$ 4 and 5, N=N=8 and 16, M=M=2 and 4 are given, respectively. As depicted in Fig. 3, BER decreases monotonically with an increase in $N$. This trend underscores the significant impact of the RIS in enhancing BER performance. For the case with $N_{r}=4$, $N=16$, and $M=2$, we can find that the BER of the RASM scheme is lower than $10^{-5}$ when SNR is equal to 10 dB and bpcu is equal to 4, which indicates the proposed scheme can achieve the desired BER performance. Particularly as the SNR increases. the simulation results closely approach to the theoretical ones, which is mainly because of the applications of CLT in theoretical analysis. Additionally, the gap between the theoretical and simulation results decreases as $N$ increases. This is also because the CLT utilization, causing the theoretical BER to converge with the simulations as $N$ grows. Moreover, we can observe that the BER performance worsens as $N_{r}$ increases, which is because the increase in $N_{r}$ leads to the increase in number of ACs, thus leading to increase error possibility in detection. For the same reason, by comparing the results of $M=2$ and $M=4$ between (a) and (b) in Fig. 3, the slight discrepancy and learn that the BER performance deteriorates with the increase of $M$. Fig. 3 (c) illustrates the analytical and simulation BER performance of the RASSK scheme, where $N_{r}=$3, 5, 6 and N=N=8, 16, respectively. Similar to (a) and (b), we can learn from (c) that BER performance of the RASSK scheme gets better with the increase of SNR and the analytical BER is approached to the simulation one especially in the higher SNR range. Additionally, the results in (c) also point out that the RASSK has satisfactory BER performance even under the low SNR situation.
[…]
Giving the same condition on SE, the comparison of the various received RIS-assisted IM schemes with the imperfect CSI estimation is shown in Fig. 5, where the schemes include the RASM, RASSK, and RSM schemes. Here, the imperfect CSI is assumed in $h_{2}^{i,m}$ with $\delta^{2}$ in the RTSM and RTSSK schemes representing the variance of estimation error [31]-[32] and the spectral efficiency is R=R=4 bpcu. Detail explanations of the imperfect CSI estimation can be found in Appendix B. Results in Fig. 5 show that the robustness of the RASSK scheme is better than that of the RASM and RSM schemes, despite the fact that the $\delta^{2}$ in the RASSK results is larger than others. Also, the BER performance of RASM scheme outperforms the one of the RSM scheme. According to the comparisons in Fig. 5, we can find that the proposed RASSK scheme has more potential to apply in complex wireless environment, due to its better robustness.
Fig. 6 shows the BER performance of RASM scheme with $N_{r}=8$$N_{d}=4$,$N=32$, and $M=2$, RASSK scheme with $N_{r}=8$ and $N_{d}=5$, RSM scheme with $N_{r}=8$, $N=32$, and $M=2$, of which the K=K=5, 10, and 15, respectively. In Fig. 6, we can find that the BER performance of all schemes gets worse with the increase of $K$. This is because stronger LoS component makes the channel responses at different antennas more similar. When receive antennas are closely spaced, their channel characteristics become very similar, thus reducing spatial diversity. Since IM schemes, especially the SM and SSK related schemes, rely on distinct channel characteristics for correct detection of antenna indices, this reduced diversity can lead to degrade the BER performance. This phenomenon is particularly evident in the results of RASSK scheme shown in Fig.6, which further corroborates the aforementioned argument. However, the BER performance of RASM scheme still outperforms the one of RSM scheme even with different $K$.
Figure 5: BER performance of the RASM, RASSK, and RASM schemes with imperfect CSI estimation.
Figure 6: BER performance of the RASM, RASSK, and RASM schemes with Rician channel fading in $\mathbf{H}_{2}$.
Fig. 7 shows the SR of the RASM and RASSK schemes. First, Fig. 7 shows that the SR of both schemes gets better with the increase of the SNR. Nevertheless, we can also observe that the increase rate stabilizes across all results within the high SNR regime. With the $N$ increasing, the SR of the RASM scheme with $N_{r}=3$, $N=8$, and $M=2$ is slightly worse than that with $N=16$, which shows the increase of $N$ can improve the PLS performance due to the enhancement in SNR at receiver. In addition, by comparing the results of the RASM scheme with different bpcu, i.e., $N_{r}=$3, 4 and M=M=2, 4, the SR increases while the bpcu increasing, thus the higher SE provides better PLS performance in the RASM scheme. Specifically, the SR of the RASM scheme with $N_{r}=3$, $N=16$, and $M=2$ is worse than that of the one with $N_{r}=4$, $N=16$, and $M=4$, which not only can be learned from the theoretical analysis in Sec. IV where the $R_{E}$ decreases with the increase in $M$ due to the $\frac{1}{M}$ term in Eq. (76), but also demonstrates the higher-order modulation is more capable of enhancing the data rate with the assistance of the RIS than the ACs indices. For the RASSK scheme, since it only transmits bits mapped from the ACs selection, the entire signal cannot be successfully detected by Eve. Thus, the SR of the RASSK is higher than that of the RASM scheme, which can be obtained by comparing the results of the RASSK scheme and the RASM scheme of the Fig. 7 (a) and (b), respectively. As also given in Fig. 7, it is clear that the SR of the RASSK scheme also increases with the rise of SNR and $N_{r}$. In contrast to the RASM scheme, the increase rate of SR in the RASSK scheme starts to level off when SNR<<0 dB, which also can indicate the superiority in PLS of the RASSK scheme.
Following model of MRC in [33], the SR of the RIS-assisted SIMO scheme with MRC (SIMO-MRC), the TSM scheme with MRC (TSM-MRC), the TSM-MRC with AN  [^9] This AN is designed with the noise variance $N_{0}$ as the same to the Gaussian white noise. Detail AN model can be found in [34], which is not presented in this paper., and the TSSK scheme with MRC (TSSK-MRC) are given to compare with the one in both proposed schemes as shown in Fig. 7.101010Since there is currently no existing literature discussing and investigating the PLS performance of the TASM scheme, where the adaptive antenna selection works at transmitter and the RIS is employed as a passive relay, we thus select the TSM and TSSK schemes as benchmarks to compare the SR with both proposed schemes in this paper. To make a fair comparison, as the same SE in 3 bits/s/Hz, the RASM has $N_{r}=$4, N=N=8, M=M=2, and the RASSK has $N_{r}=$4, N=N=8, and the SIMO-MRC has $N_{t}=$1 $N_{r}=$3, N=N=8, M=M=32, and the TSM-MRC scheme has $N_{t}=$3 $N_{r}=$3, N=N=8, M=M=2, and the TSSK-MRC scheme has $N_{t}=$8 $N_{r}=$3, N=N=8. In Fig. 7, the SR of both proposed schemes is higher than that of the others from -30 dB to 30 dB, demonstrating that the received IM schemes achieve better PLS performance compared to the traditional SIMO-MRC, TSM-MRC, and TSSK-MRC schemes, even the PLS wireless technique, e.g., AN in the TSM-MRC. Below around -5 dB, the SR of the SIMO-MRC scheme is slightly higher than that of TSM-MRC and TSSK-MRC, indicating that the PLS performance based on transmit antenna indices is weaker than that based on constellation symbols. The SR results of TSM-MRC and TSSK-MRC are closed because Bob and Eve share the same transmitter-to-RIS channel, and the CSI is assumed to be known to Eve, allowing her to fully detect and decode the transmitted signals, which is a limitation of transmitted IM schemes. Notably, the SR of the SIMO-MRC, TSM-MRC, and TSSK-MRC schemes approaches but does not reach zero, as Bob’s channel capacity remains higher than Eve’s due to having multiple antennas, while Eve only has one.
Figure 7: (a) SR of the RASM scheme and (b) the RASSK scheme, with $N_{r}=$3, 4, and 5, N=N=8 and 16, M=M=2 and 4.
Figure 8: SR of the RASM and RASSK schemes compared to the RIS-assisted SIMO, TSM, TSSK schemes empowered by the MRC with $N_{r}=$3 and 4, $N_{t}=$1, 4, and 8, M=M=2 and 32, N=N=8, and AN.
VII Conclusion
This paper presented two novel RIS-assisted IM schemes, RASM and RASSK, aimed at enhancing SE and ensuring PLS in wireless transmissions. A comprehensive analysis of SE, BER, detection complexity, and SR confirms that the proposed schemes provided reliable communication quality and protect confidential information from eavesdropping. Compared to traditional IM schemes, they offered better SE, PLS performance, and robustness against imperfect CSI. To further approach practical applications, more sophisticated eavesdropper model will be considered in our future works. Overall, these schemes showed strong potential for future wireless applications.
Appendix A
The mutual information of Bob and Eve are derived. The achievable data rate for Bob and Eve can be obtained as a discrete-input continuous-output memoryless channel (DCMC) capacity [29]-[30]. For the RASM scheme, mutual information of Bob can be obtained as:
$\mathfrak{T}_{b}\left(k,r;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2}\right)=$ $\mathfrak{T}_{b}\left(k;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2}\right)+$
$\mathfrak{T}_{b}\left(r;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2},k\right),$ (68)
where $y_{m}$ represents the received signal at $m$-th receive antenna, $\mathfrak{T}_{b}\left(k;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2}\right)$ represents the mutual information between $\boldsymbol{Y}$ and $x_{k}$, $\mathfrak{T}_{b}\left(r;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2},k\right)$ stands for the mutual information between $\boldsymbol{Y}$ and $r$-th AC after AC selection when $x_{k}$ is given. $\mathfrak{T}_{b}\left(k;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2}\right)$ can be further expressed as:
$\mathfrak{T}_{b}\left(k;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2}\right)$
$=\sum_{k=1}^{M}{\int{P_{r}\left(x_{k}\right)f\left(\boldsymbol{Y}\middle|k\right)}\log_{2}\!\>\frac{f\left(\boldsymbol{Y}\middle|k\right)}{f\left(\boldsymbol{Y}\right)}d\boldsymbol{Y}},$ (69)
where $P_{r}\left(x_{k}\right)=\frac{1}{M}$ since each constellation symbol is selected equiprobably, which is the same in each AC equiprobably selected in $\varLambda(r)$ as $P_{r}\left(r\right)=\frac{1}{D}$. And the PDF of $\boldsymbol{Y}$ and the conditional PDF of $\boldsymbol{y}$ when $k$ is given are $f\left(\boldsymbol{Y}\right)=\sum_{r=1}^{D}{\sum_{k=1}^{M}{\frac{e^{-\!\>\left\|\boldsymbol{Y}-\mathbf{G}_{r,k}\right\|_{2}^{2}}}{DM\pi^{N_{a}}}}}$ and $f\left(\boldsymbol{Y}\middle|k\right)=\frac{1}{D}\sum_{r=1}^{D}{\frac{e^{-\!\>\left\|\boldsymbol{Y}-\mathbf{G}_{r,k}\right\|_{2}^{2}}}{D\pi^{N_{a}}}},$ respectively. By substituting $f\left(\boldsymbol{Y}\right)$ and $f\left(\boldsymbol{Y}\middle|k\right)$ into (65) and after some algebra, the mutual information between $y_{l}$ and $x_{k}$ can be rewritten as
$\mathfrak{T}_{b}\left(k;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2}\right)=\log_{2}\!\>M$-
$\frac{1}{DM}\sum_{r=1}^{D}{\sum_{k=1}^{M}{\sum_{\hat{r}=1}^{D}{\sum_{\hat{k}=1}^{M}{\mathbb{E}_{\tau_{n}}\left[\log_{2}e^{\!\>\left\|\boldsymbol{Y}-\mathbf{G}_{\hat{r},\hat{k}}\right\|_{2}^{2}-\!\>\left\|\boldsymbol{Y}-\mathbf{G}_{r,k}\right\|_{2}^{2}}\!\>\right]}}}}.$ (70)
Also, the mutual information between $y_{l}$ and $r$-th AC in $\varLambda(r)$ when $x_{k}$ is given can be obtained as:
$\mathfrak{T}_{b}$ $\left(r;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2},k\right)$
$=\sum_{r=1}^{D}{\sum_{k=1}^{M}{\int{f\left(k,r\right)f\left(\boldsymbol{y}\middle|k,r\right)\log_{2}\!\>\frac{f\left(\boldsymbol{y}\middle|k,r\right)}{f\left(\boldsymbol{y}\middle|k\right)}dy_{l}}}},$ (71)
where $f\left(k,r\right)$ represents a joint PDF of $x_{k}$ and $r$-th AC in $\varLambda(r)$ as a product of their individual PDFs which follows the uniform distribution function, and $f\left(\boldsymbol{y}\middle|k,r\right)$ denotes the conditional PDF of $\boldsymbol{y}$ when $k$ and $r$ are given as:
$f\left(\boldsymbol{Y}\middle|k,r\right)=\frac{1}{\pi^{N_{a}}}\exp\left(-{\left\|\mathbf{N}\right\|_{2}^{2}}\right).$ (72)
Substituting $f\left(\boldsymbol{Y}\middle|k\right)$ and Eq. (71) into Eq. (69), the mutual information between $y_{l}$ and $r$-th AC in $\varLambda(r)$ is:
$\mathfrak{T}_{b}\left(r;\boldsymbol{Y}\middle|\mathbf{H}_{1},\mathbf{H}_{2},k\right)=\log_{2}\!\>D$-
$\frac{1}{DM}\sum_{r=1}^{D}{\sum_{k=1}^{M}{\mathbb{E}_{\tau_{n}}\left[\log_{2}\!\>\sum_{r=1}^{D}{e^{-\left\|\boldsymbol{Y}-\mathbf{G}_{r,k}\right\|_{2}^{2}+\mathbf{N}}}\right]}}.\!\>$ (73)
By taking expectation of Eq. (73) and Eq. (75) over $\mathrm{h}_{1}^{i}$ and $\mathrm{h}_{2}^{i,n}$, $R_{B}^{r}$ and $R_{B}^{k}$ can be obtained respectively. Finally, according to Eq. (62), the data rate of Bob can be written as Eq. (64). Since Eve only can detect the constellation symbols of the transmitted signals, we can consider that $R_{E}^{r}=0$, thus the Eq. (63) can be further rewritten as $R_{E}=\frac{1}{M}R_{E}^{k}$. Besides, assuming both Eve and Bob can perfectly detect the constellation symbols, the mutual information of Eve is similar to $\mathfrak{T}_{b}\left(k;y_{l}\middle|\mathrm{h}_{1}^{i},\mathrm{h}_{2}^{i,l}\right)$ as given in Eq. (69), which can be derived as follows:
$\mathfrak{T}_{e}\left(k;y_{e}\middle|\mathrm{g}_{1}^{i},\mathrm{g}_{2}^{i}\right)=\sum_{k=1}^{M}{\int{P_{r}\left(x_{k}\right)f\left(y_{e}\middle|k\right)}\log_{2}\!\>\frac{f\left(y_{e}\middle|k\right)}{f\left(y_{e}\right)}dy_{e}},$ (74)
with $f\left(y_{e}\middle|k\right)=\frac{1}{\pi}\exp\left(-\left\|\tau_{e}\right\|_{2}^{2}\right)$ and
$f\left(y_{e}\right)=\frac{1}{M}\sum_{k=1}^{M}{\frac{1}{\pi}\exp\left(-\left\|y_{e}-\sum_{i=1}^{N}{g_{2}^{i}\Phi_{i}h_{1}^{i}x_{k}}\right\|_{2}^{2}\right)}.$ (75)
Then, by taking expectation of Eq. (75) over $g_{1}^{i}$ and $g_{2}^{i}$ and multiplying $\frac{1}{M}$, we can obtain the achievable data rate of Eve as Eq. (65).
Appendix B
The CSI estimation at receiver can hardly be perfect especially in complex wireless environment. For simplicity, the imperfect CSI estimation is only considered in the channel from the RIS to receive. According to this and referring to the imperfect CSI model from [31]-[32], the coefficient of $\check{h}_{2}^{i,m}$ can be expressed as the sum of $\dot{h}_{2}^{i,m}\sim\mathcal{C}\mathcal{N}(0,1-\delta_{e1}^{2})$ and $\ddot{h}_{2}^{i,m}\sim\mathcal{C}\mathcal{N}(0,\mathrm{}\delta_{e1}^{2})$. Giving that $\delta_{e2}^{2}$ denotes the inaccuracy of estimation in CSI with $\ddot{\delta_{e2}^{2}}=1-\delta_{e2}^{2}$. Thus, we can respectively obtain the receive signals at $l$-th receive antenna in $r$-th selected AC for the RASM and RASSK schemes with imperfect CSI estimation as
$\ddot{y}_{l}^{RASM}=\underset{\mathrm{constructive}\,\,part}{\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{N_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}x_{k}}}+$
$\underset{\mathrm{non}-\mathrm{constructive}\,\,part}{\underbrace{\sum_{q=1,\mathrm{ }q\neq l}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}x_{k}}}}+\tau_{l},$ (76)
$\ddot{y}_{l}^{RASSK}=\underset{\mathrm{constructive}\,\,part}{\sqrt{E_{s}}\underbrace{\sum_{i=\left(l-1\right)N_{E}+1}^{N_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}}}+$
$\sqrt{E_{s}}\underset{\mathrm{non}-\mathrm{constructive}\,\,part}{\underbrace{\sum_{q=1,\mathrm{ }q\neq l}^{N_{a}}{\sum_{i=\left(q-1\right)N_{E}+1}^{qN_{E}}{\left(\delta_{e2}^{2}\dot{h}_{2}^{i,m}+\ddot{\delta}_{e2}^{2}\ddot{h}_{2}^{i,m}\right)\Phi_{i,l}h_{1}^{i}}}}}+\tau_{l}.$ (77)